Swan Company has a direct labor standard of 15 hours per unit of output. Each employee has a standard wage of $14 per hour. During March, employees worked 13,100 hours. The direct labor rate variance was $9,170 favorable, the direct labor efficiency variance was $15,400 unfavorable. What was the actual payroll?
Answer #1

Calculation of Actual Rate using Labor rate variance :-

Labor rate variance = (Actual rate - standard rate) x actual hours

9,170 F = (Actual rate - 14) x 13,100

-9,170/13,100 = Actual rate - 14

-0.70 = actual rate - 14

Actual rate = 14 - 0.70 = $13.30

Actual labor rate per hour = $13.30

Actual Payroll = actual hours x actual rate

= 13,100 x 13.30 = $174,230